# Linear response and exact hydrodynamic projections in Lindblad equations with decoupled Bogoliubov hierarchies

### v1

@@ IV.1.1 Eigenvalue spectrum for Model I
Figure 1: Complex eigenvalues as a function of total momentum for Model I with $\gamma=3$, $J=1$ and $L=100$. The real parts are approximately $-2\gamma$, indicating that the corresponding modes are strongly damped.
@@
We see that there is a continuum of eigenvalues with non-vanishing imaginary parts and real parts close to $-2\gamma$, cf. Fig. 1. In addition there is a band of real eigenvalues that tends to zero as $(p_{n}-\pi)^{2}$ for total momentum $p_{n}\approx\pi$. This reflects the existence of a diffusive hydrodynamic mode related to the particle number $U(1)$ symmetry of the model. The mode occurs around $p_{n}=\pi$ in our two-particle problem because of the alternating sign introduced in the vectorization (13) of the annihilation operators $c_{j}$. This leads to a shift in total momentum by $\pi$ in the vectorized formalism compared to the underlying equation of motion for the operators ${\cal O}(\vec{j},\vec{\ell})$.
Both types of solutions shown in Figs 1 and 2 are easily understood in the infinite volume limit, in which both $p_{n}$ and $q_{n}$ turn into continuous variables. The two classes of solutions shown in Figs 1 and 2 respectively then give rise to eigenvalues of the following form:
@@
  • $q=\kappa+i\eta\in\mathbb{C}$ with $\kappa=-\frac{\pi}{2}{\rm sgn}\left(J\cos(\frac{p}{2})\right)$ and $\eta=\ \mathrm{arccosh}\Big{(}\big{|}\frac{\gamma}{2J\cos(p/2)}\big{|}\Big{)}$. This gives rise to a single momentum-dependent mode with real eigenvalues
@@
The wave-function (46) describes a two-particle bound state, and is the analog [21] of the $k$-$\Lambda$-string excitation in the one-dimensional Hubbard model [61, 57] in the case where the hopping amplitude is purely imaginary.
IV.1.2 Eigenvalue spectrum for Model II
The same two classes of solutions as in Model I exists here as well. In the $L\to\infty$ limit solutions with real $k_{1,2}$ give rise to the same two-parameter continuum of states as in Model I, cf. Fig. 1. Complex solutions to the quantization conditions (39) have the form $q=\kappa+i\eta$ with $\kappa=\frac{\pi}{2}{\rm sgn}\left(J\cos(\frac{p}{2})\cos(p)\right)$ and $\eta~{}=~{}\mathrm{arccosh}\left(\left|\frac{\gamma\cos(p)}{2J\cos(p/2)}\right|\right)$ and correspond to momentum dependent real eigenvalues of the form
@@ IV.1.2 Eigenvalue spectrum for Model II
The wave function corresponding to this state in the large-$L$ limit can again be written in the form (46).
@@
IV.1.3 Eigenvalue spectrum for Model III
@@
IV.1.4 Eigenvalue spectrum for Model IV
@@
IV.2 Position representation for the eigenvalue equation with $n_{\uparrow}=2$
@@
The vectorized form of their equations of motion is given by (14) with $n_{\downarrow}=0$, $n_{\uparrow}=2$. To proceed we again solve the corresponding eigenvalue equation (III.1) in the position representation, which reads
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\sigma=\pm 1}}\left(\delta_{x,x^{\prime}+\sigma}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\sigma}\right)$
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\sigma=\pm 1}}(\delta_{x,x^{\prime}+\sigma}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\sigma})$
$-\gamma\Big{(}2+\beta\sum_{\sigma=\pm 1}\delta_{x,y+\sigma}\Big{)}\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}\ .$ (60)
@@
IV.2.1 Eigenvalue spectrum for Model I
@@
Figure 8: Eigenvalues for $n_{\uparrow}=2$ as a function of total momentum for Model I with $\gamma=3$, $J=1$ and $L=100$. The real parts are $-2\gamma$ leading to strongly damping of the corresponding modes.
We see that all eigenvalues have a real part of $-2\gamma$. This implies that all operators ${\cal O}(x,y)$ decay exponentially in time. This is expected for charged operators.
IV.2.2 Eigenvalue spectra for Models II, III and IV
@@
V Hydrodynamic projections
We have seen that in the models with U(1) symmetry there exists a single-particle mode in the $n_{\uparrow}=n_{\downarrow}=1$ sector, which has a vanishing real part at $p=\pi$. Clearly this mode will dominate the late-time behaviour. The corresponding eigenstates are
@@ V Hydrodynamic projections
where the wave functions are given by (46). At late times we may therefore simplify the expression (21) for the time evolved state by restricting the sum over eigenstates to the subset of states (69)
@@ A. Operators carrying a definite momentum: As an example we consider
This shows that operators carrying a definite momentum do not acquire hydrodynamic power-law (in $t$) tails.
@@
$\psi_{x_{0},y_{0}}(x,y;t)$ $\simeq(-1)^{y+y_{0}}\int_{0}^{2\pi}\frac{\mathrm{d}p}{2\pi}\coth(\eta(p))e^{\lambda(p)t+i\frac{p}{2}Y}$
×$\qquad\times\ e^{(i\kappa(p)-\eta(p))X}\ ,$ (82)
@@
$\kappa(p)$ $=-\frac{\pi}{2}{\rm sgn}\big{(}\cos\big{(}\frac{p}{2}\big{)}\big{)}\ .$ (83)
@@
$\psi_{x_{0},y_{0}}(x,y;t)\simeq(-1)^{y+y_{0}}\int_{0}^{\pi}\frac{\mathrm{d}p}{2\pi}\Big{[}\frac{J(\pi-p)}{2\gamma}\Big{]}^{X}e^{-\alpha t(\pi-p)^{2}}$
×$\times c(\pi-p)\left[e^{i\frac{p}{2}Y-i\frac{\pi}{2}X}+(-1)^{Y}e^{-i\frac{p}{2}Y+i\frac{\pi}{2}X}\right]\ .$ (86)
@@
$\psi_{x_{0},y_{0}}(x,y;t)$ $\approx(-1)^{\min(x-y,0)+\max(x_{0}-y_{0},0)}\frac{1}{2\pi}\big{(}\frac{J}{2\gamma}\big{)}^{X}\Big{[}g(X,\alpha t,Y)+\frac{1}{2}\frac{\mathrm{d^{2}}c(q)}{\mathrm{d}q^{2}}\bigg{|}_{q=0}g(X+2,\alpha t,Y)+\dots\Big{]},$
$g(n,\alpha,Y)$ $=\frac{n!}{2^{n}}(\alpha t)^{-\frac{1+n}{2}}\begin{cases}\text{Re}\left(U(\frac{1+n}{2},\frac{1}{2},-\frac{Y^{2}}{16\alpha t})\right)&\text{if }Y\mathrm{even},\\ i\ {\rm sgn}{(Y)}\ \text{Im}\left(U(\frac{1+n}{2},\frac{1}{2},-\frac{Y^{2}}{16\alpha t})\right)&\text{if }Y\mathrm{odd}.\end{cases}$ (88)
@@
$\psi_{x_{0},y_{0}}(x,y;t)\Bigg{|}_{\rm leading}=\begin{cases}A_{e}\ (\alpha t)^{-\frac{1+X}{2}}&\text{Y even}\ ,\\ A_{o}\ (\alpha t)^{-\frac{2+X}{2}}&\text{Y odd}\ ,\end{cases}$ (89)
@@
To summarize, we have established that at late times $\gamma t\gg 1$ the Heisenberg-picture operator takes the following form
@@
This operator spreads diffusively, as is shown in Fig. 11.
@@
V.1 Hydrodynamic projections in Models II, III, IV
Hydrodynamic projections in Models II, III and IV can be determined along the same lines as in Model I. The wave function of the hydrodynamic mode has the same form in terms of the parameters $\kappa_{n}$ and $\eta_{n}$, and all we have to do is to substitute the appropriate expressions for these in Models II-IV. A. Operators carrying a definite momentum: The operators $\hat{\Psi}_{1}^{q}$ defined in (72) decay exponentially in time. The decay rate $\Delta(q)$ is given in terms of the eigenvalues $\lambda(p)$ (48) (Model II),(49), (52) (Model III) and (53), (56) (Model IV) as $\Delta(q)=\max(\mathrm{Re}(\lambda(q+\pi)))$. Here the maximum is taken over the different modes at fixed momentum. In Model II a simple closed-form expression is available
@@ V.1 Hydrodynamic projections in Models II, III, IV
$\alpha$ $=\frac{(J^{2}+\gamma^{2})}{\gamma}\ .$ (95)
The same analysis for Models III and IV leads again to (88), where $\alpha$ and $c(q)$ are now given by
@@
$\alpha$ $=\frac{2J^{2}+(2+\xi)\gamma^{2}}{(2+\xi)\gamma}\ .$ (97)
@@
VI Dynamics of operators quadratic in fermions in Model V
We now turn to Model V, which does not have particle number conservation and we therefore do not expect any hydrodynamic behaviour at late times. On the other hand, the absence of particle number conservation makes the solution of the equations of motion in the Heisenberg picture more interesting. The vectorized Hamiltonian (19) has the key property that its interaction terms in ${\cal H}$ contain more annihilation that creation operators, which results in a triangular structure of the corresponding eigenvalue equations, which facilitates their solution.
VI.1 Position representation for the eigenvalue equation with at most $n_{\uparrow}=n_{\downarrow}=1$ fermions
@@ VI Dynamics of operators quadratic in fermions in Model V
The eigenvalue equations of the Hamiltonian ${\cal H}$ needed to work out $|\Psi_{1,1}(t)\rangle$ are
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\sigma=\pm 1}}\left(\delta_{x,x^{\prime}+\sigma}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\sigma}\right)$
$-2\gamma\Big{(}1-\sum_{\sigma=\pm 1}\delta_{x,y+\sigma}\Big{)}\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}$
$+\delta_{x,y}\sum_{\sigma=\pm 1}\delta_{x+\sigma,x^{\prime}}\delta_{y+\sigma,y^{\prime}}\ .$ (104)
@@
Substituting (110) into the eigenvalue equation (103) imposes that either $v_{\alpha}=0$ or $\lambda_{\alpha}=-4\gamma$. We find numerically that the eigenvalue $-4\gamma$ is non-degenerate. The corresponding eigenoperator is given above, which establishes our assertion.
VI.1.1 Eigenvalue spectrum

### v2

@@ V.1.1 Eigenvalue spectrum for Model I
Figure 1: Complex eigenvalues as a function of total momentum for Model I with $\gamma=3$, $J=1$ and $L=100$. The real parts are approximately $2\gamma$, indicating that the corresponding modes are strongly damped.
@@
We see that there is a continuum of eigenvalues with non-vanishing imaginary parts and real parts close to $2\gamma$, cf. Fig. 1. In addition there is a band of real eigenvalues that tends to zero as $(p_{n}-\pi)^{2}$ for total momentum $p_{n}\approx\pi$. The mode occurs around $p_{n}=\pi$ in our two-particle problem because of the alternating sign introduced in the vectorization (26) of the annihilation operators $c_{j}$. This leads to a shift in total momentum by $\pi$ in the vectorized formalism compared to the underlying equation of motion for the operators ${\cal O}(\vec{j},\vec{\ell})$. In the original (unvectorized) problem the mode therefore corresponds to a family of eigenoperators of the time evolution of the form
$\hat{\Phi}(P)=\sum_{x,y}\Phi_{R,P+\pi}(x,y)(-1)^{y}c^{\dagger}_{x}c_{y}\ ,$ (63)
which have eigenvalues
$\lambda(P\rightarrow 0)=-\frac{J^{2}}{\gamma}P^{2}+\dots.$ (64)
For small momenta these are diffusive hydrodynamic modes related to the particle number $U(1)$ symmetry of the model, see the discussion in the introduction. Both types of solutions shown in Figs 1 and 2 are easily understood in the infinite volume limit, in which both $p_{n}$ and $q_{n}$ turn into continuous variables. The two classes of solutions shown in Figs 1 and 2 respectively then give rise to eigenvalues of the following form:
@@
  • $q=\kappa+i\eta\in\mathbb{C}$ with $\kappa=-\frac{\pi}{2}{\rm sgn}\left(J\cos(\frac{p}{2})\right)$ and $\eta=\ \mathrm{arccosh}\Big(\big|\frac{\gamma}{2J\cos(p/2)}\big|\Big)$. This solution exists if $|\frac{2J\cos(p/2)}{\gamma}|<1$ and gives rise to a single momentum-dependent mode with real eigenvalues
@@
The wave-function (67) describes a two-particle bound state, and is the analog [51] of the $k$-$\Lambda$-string excitation in the one-dimensional Hubbard model [64, 25] in the case where the hopping amplitude is purely imaginary. The fact that we are dealing with a bound state solution implies that the associated eigenoperators (19) of the Lindbladian are sums of terms that are exponentially localized around the centre of mass between the creation and annihilation operator.
V.1.2 Eigenvalue spectrum for Model II
The same two classes of solutions as in Model I exists here as well. In the $L\to\infty$ limit solutions with real $k_{1,2}$ give rise to the same two-parameter continuum of states as in Model I, cf. Fig. 1. Complex solutions to the quantization conditions (58) have the form $q=\kappa+i\eta$ with $\kappa=\frac{\pi}{2}{\rm sgn}\left(J\cos(\frac{p}{2})\cos(p)\right)$ and $\eta~=~\mathrm{arccosh}\left(\left|\frac{\gamma\cos(p)}{2J\cos(p/2)}\right|\right)$ and exists provided that $|\frac{2J\cos(p/2)}{\gamma\cos(p)}|<1$. They correspond to momentum-dependent real eigenvalues of the form
@@ V.1.2 Eigenvalue spectrum for Model II
The dispersion (69) is shown in Fig. 3 for $\gamma=3$ and $L=100$. The wave function corresponding to this state in the large-$L$ limit can again be written in the form (67).
@@
V.1.3 Eigenvalue spectrum for Model III
@@
V.1.4 Eigenvalue spectrum for Model IV
@@
V.2 Position representation for the eigenvalue equation with $N_{\uparrow}=2$
@@
The vectorized form of their equations of motion is given by (27) with $N_{\downarrow}=0$, $N_{\uparrow}=2$. To proceed we again solve the corresponding eigenvalue equation (33) in the position representation, which reads
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\tau=\pm 1}}\left(\delta_{x,x^{\prime}+\tau}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\tau}\right)$
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\tau=\pm 1}}(\delta_{x,x^{\prime}+\tau}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\tau})$
$-\gamma\Big(2+\beta\sum_{\tau=\pm 1}\delta_{x,y+\tau}\Big)\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}\ .$ (81)
@@
V.2.1 Eigenvalue spectrum for Model I
@@
Figure 8: Eigenvalues for $N_{\uparrow}=2$ as a function of total momentum for Model I with $\gamma=3$, $J=1$ and $L=100$. The real parts are $2\gamma$ leading to strongly damping of the corresponding modes.
We see that all eigenvalues have a real part of $2\gamma$. This implies that all operators ${\cal O}(x,y)$ decay exponentially in time. This is expected for charged operators.
V.2.2 Eigenvalue spectra for Models II, III and IV
@@
VI Hydrodynamic projections
We now turn to the problem of hydrodynamic projections in the $N_{\uparrow}=N_{\downarrow}=1$ sector. We have seen that in the models with U(1) symmetry there exists a single-particle mode in this sector, which has a vanishing real part at $p=\pi$. Clearly this mode will dominate the late-time behaviour. The corresponding eigenstates are
@@ VI Hydrodynamic projections
where the wave functions are given by (67). At late times we may therefore simplify the expression (34) for the time-evolved state by restricting the sum over eigenstates to the subset of states (90)
@@ A. Operators carrying a definite momentum: As an example we consider
For $(q,d)=(\frac{\pi}{5},10)$ we observe two regimes: at early times the operator decays exponentially with a rate of approximately $2\gamma$, which is set by the real part of the two-particle continuum shown in Fig. 1. At late times the decay rate is set by the bound state solution shown in Fig. 2. As we are considering an operator with definite momentum, only a single momentum of the bound state solution will contribute (as long as the latter exists). However, this has a very small overlap with the operator under consideration, which is why this behaviour only becomes visible at late times, when the contributions due to the two-particle continuum have already decayed to negligible values.
Our results show that quadratic operators carrying a definite momentum do not acquire diffusive power-law (in $t$) tails. This is a result of the kinematics of the hydrodynamic modes, which is a direct consequence of the decoupling of the BBGKY hierarchy.
@@
$\psi_{x_{0},y_{0}}(x,y;t)$ $\simeq(-1)^{y+y_{0}}\int_{p_{0}}^{2\pi-p_{0}}\frac{\mathrm{d}p}{2\pi}\coth(\eta(p))e^{\lambda(p)t+i\frac{p}{2}Y}$
$\qquad\times\ e^{(i\kappa(p)-\eta(p))X}\ ,$ (103)
@@
$\kappa(p)$ $=-\frac{\pi}{2}{\rm sgn}\big(\cos\big(\frac{p}{2}\big)\big)\ ,$
$p_{0}$ $=2\arccos(\frac{\gamma}{2J})\ \theta(2J-\gamma)$ (104)
@@
$\psi_{x_{0},y_{0}}(x,y;t)\simeq(-1)^{y+y_{0}}\int_{p_{0}}^{\pi}\frac{\mathrm{d}p}{2\pi}\Big[\frac{J(\pi-p)}{2\gamma}\Big]^{X}e^{-Dt(\pi-p)^{2}}$
$\times c(\pi-p)\left[e^{i\frac{p}{2}Y-i\frac{\pi}{2}X}+(-1)^{Y}e^{-i\frac{p}{2}Y+i\frac{\pi}{2}X}\right]\ .$ (107)
@@
$\psi_{x_{0},y_{0}}(x,y;t)$ $\approx(-1)^{\min(x-y,0)+\max(x_{0}-y_{0},0)}\frac{1}{2\pi}\big(\frac{J}{2\gamma}\big)^{X}\Big[g(X,Dt,Y)+\frac{1}{2}\frac{\mathrm{d^{2}}c(q)}{\mathrm{d}q^{2}}\bigg|_{q=0}g(X+2,Dt,Y)+\dots\Big],$
$g(n,Dt,Y)$ $=\frac{n!}{2^{n}}(Dt)^{-\frac{1+n}{2}}\begin{cases}\text{Re}\left(U(\frac{1+n}{2},\frac{1}{2},-\frac{Y^{2}}{16Dt})\right)&\text{if }Y\mathrm{even},\\ i\ {\rm sgn}{(Y)}\ \text{Im}\left(U(\frac{1+n}{2},\frac{1}{2},-\frac{Y^{2}}{16Dt})\right)&\text{if }Y\mathrm{odd}.\end{cases}$ (109)
@@
$\psi_{x_{0},y_{0}}(x,y;t)\Bigg|_{\rm leading}=\begin{cases}A_{e}\ (Dt)^{-\frac{1+X}{2}}&\text{Y even}\ ,\\ A_{o}\ (Dt)^{-\frac{2+X}{2}}&\text{Y odd}\ ,\end{cases}$ (110)
@@
Some of these power-laws were recently calculated in [5] using a transfer matrix approach. To summarize, we have established that at late times $\gamma t\gg 1$ the Heisenberg-picture operator takes the following form
@@
This operator spreads diffusively, as is shown in Fig. 11 (see also [14]).
@@
VI.1 Hydrodynamic projections in Models II, III, IV
Hydrodynamic projections in Models II, III and IV can be determined along the same lines as in Model I. The wave function of the hydrodynamic mode has the same form in terms of the parameters $\kappa_{n}$ and $\eta_{n}$, and all we have to do is to substitute the appropriate expressions for these in Models II-IV.
A. Operators carrying a definite momentum: The operators $\hat{\Psi}_{1}^{q}$ defined in (93) decay exponentially in time. The decay rate $\Delta(q)$ is given in terms of the eigenvalues $\lambda(p)$ (69) (Model II),(70), (73) (Model III) and (74), (77) (Model IV) as $\Delta(q)=\max(\mathrm{Re}(\lambda(q+\pi)))$. Here the maximum is taken over the different modes at fixed momentum. In Model II a simple closed-form expression is available
@@ VI.1 Hydrodynamic projections in Models II, III, IV
$D$ $=\frac{(J^{2}+\gamma^{2})}{\gamma}\ .$ (116)
The same analysis for Models III and IV leads again to (109), where $D$ and $c(q)$ are now given by
@@
$D$ $=\frac{2J^{2}+(2+\xi)\gamma^{2}}{(2+\xi)\gamma}\ .$ (118)
@@
VII Dynamics of operators quadratic in fermions in Model V
[…]
VII.1 Position representation for the eigenvalue equation with at most $N_{\uparrow}=N_{\downarrow}=1$ fermions
@@ VII Dynamics of operators quadratic in fermions in Model V
[…]
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\tau=\pm 1}}\left(\delta_{x,x^{\prime}+\tau}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\tau}\right)$
$-2\gamma\Big(1-\sum_{\tau=\pm 1}\delta_{x,y+\tau}\Big)\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}$
$+\delta_{x,y}\sum_{\tau=\pm 1}\delta_{x+\tau,x^{\prime}}\delta_{y+\tau,y^{\prime}}\ .$ (125)
@@
Substituting (131) into the eigenvalue equation (124) imposes that either $v_{\alpha}=0$ or $\lambda_{\alpha}=-4\gamma$. We find numerically that the eigenvalue $4\gamma$ is non-degenerate. The corresponding eigenoperator is given above, which establishes our assertion.
VII.1.1 Eigenvalue spectrum